OpenStudy (anonymous):

use rational exponents to simplify little 5 on the outside sqrt x 25

OpenStudy (saifoo.khan):

\[\Huge \sqrt[5]{25}\]Something like that?

OpenStudy (anonymous):

almost but inside is x^25

OpenStudy (saifoo.khan):

\[\Huge \sqrt[5]{x^{25}}\]

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[x^\frac{25}{5} = \Huge x^5\]

OpenStudy (anonymous):

Thank you

OpenStudy (saifoo.khan):

You welcome. can i ask you a question?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

nvm. do u have any more problems?

OpenStudy (anonymous):

yes but I just do not understand rational exponets to save my life. sorry my comp froze

OpenStudy (saifoo.khan):

no problem. where are u stuck? i can help you through.

OpenStudy (anonymous):

\[\sqrt[3]{125x^14/y}\]

OpenStudy (saifoo.khan):

can u draw it?

OpenStudy (saifoo.khan):

are u drawing?

OpenStudy (saifoo.khan):

i can't get the position of y.

OpenStudy (anonymous):

trying

OpenStudy (anonymous):

\[\sqrt[3]{125x^14/y^3}\]

OpenStudy (saifoo.khan):

or if you want, u can post a photo as well.

OpenStudy (anonymous):

Does that help it is a fraction on the inside

OpenStudy (anonymous):

the in side is 125x^14/ y^3